# Exploring Neuron Interactions and Emergence in LLMs: From the Multifractal Analysis Perspective

### v4

@@ 4.2 Neuron-based Multifractal Analysis (NeuroMFA)
where $D$ denotes the fractal dimension characterizing the fractal properties of the interactions among a neuron and its neighbors.
To capture the multifractality of the neuron interactions, we develop the NeuroMFA framework. This method integrates the fractal features of different neurons in the NIN. The distortion factor $q$ is introduced to differentiate the nuances of fractal structural features. For a given neuron $v_{i}$ at layer $l$ in the neural network $G=(V,E)$, the mass probability measure $p_{l,i}(r)$ at a radius $r$ is defined as:
$p_{l,i}(r)=\frac{N_{l,i}(r)}{T_{l,i}}$ (4)
where $N_{l,i}(r)$ is the number of neurons within the radius $r$ from neuron $v_{i}$, and $T_{l,i}$ is the total number of neurons in layer $l+1$ that are connected to neuron $v_{i}$.
Definition 4.3 (Partition Function).
The partition function $Z_{q}(r)$ for the neuron-based multifractal analysis is defined as the sum of the $q$-th power of the probability measures across all neurons in the network, given by:
$Z_{q}(r)=\sum_{l=1}^{L}\sum_{v_{i}\in L_{l}}p_{l,i}(r)^{q}$ (5)
We provide the algorithm to obtain the partition function in Appendix Algorithm 5.
Observation 4.2.
A log-log relationship exists between $Z_{q}(r)$ and the observational scale $r/d$, validated in Table 2, Section C.1.4:
$Z_{q}(r)\sim\left(\frac{r}{d}\right)^{\tau(q)},$ (6)
where $d$ is the maximum box radius and $\tau(q)$ is the mass exponent.
@@
then the network is said to possess multifractal scale invariance. Here, $B_{i}(\epsilon)$ represents the box of scale $\epsilon$, and $\tau(q)$ is the mass exponent.
Algorithm 1 Calculation of the Mass Exponent $\tau(q)$ in Eqn. 6
0:  $r_{\text{sorted}}$ (a list of sorted box radius, output from Algorithm 5), $Zq_{\text{list}}$ (a list of Partition functions, output from Algorithm 5), $d$ (maximum of box radius)
1:  for all $Zq$ in $Zq_{\text{list}}$ do
2:      Calculate the slope of the linear regression of $\log(Zq)$ on $\log\left(\frac{r_{\text{sorted}}}{d}\right)$
3:     Save the slopes to an empty list: $\tau_{\text{list}}$
4:  end for
5:  return $\tau_{\text{list}}$
Definition 4.4 (Multifractal Spectrum).
The Legendre transform enables the computation of the the Lipschitz-Hölder exponent $\alpha(q)$ and multifractal spectrum $f(\alpha)$ characterizing the multifractal structural features of the network:
@@
$f(\alpha)=q\alpha(q)-\tau(q),$ (9)
We provide the algorithm to obtain the partition function in Algorithm 2.
Algorithm 2 Calculation of the Multifractal Spectrum $f(\alpha)$ in Eqn. 9
0:  $\tau_{\text{list}}$ (a list of mass exponents, output from Algorithm 1), $q_{\text{list}}$ (a list of distortion factors)
1:   Calculate the $\alpha_{\text{list}}$ using Eqn.8 $\alpha_{\text{list}}\leftarrow\frac{d\tau_{\text{list}}}{dq_{\text{list}}}$
2:   Calculate the $f(\alpha)_{\text{list}}$ using Eqn.9 $f(\alpha)_{\text{list}}\leftarrow q_{\text{list}}\cdot\alpha_{\text{list}}-\tau_{\text{list}}$
3:  $\alpha_{0}\leftarrow$ $\alpha_{\text{list}}$ [index of maximum value in $f(\alpha)_{\text{list}}$]
4:  $\text{width}\leftarrow\max(\alpha_{\text{list}})-\min(\alpha_{\text{list}})$
5:  return $(\alpha_{0},\text{width},\alpha_{\text{list}},f(\alpha)_{\text{list}})$
Figure 5: The distributions of average weighted degree for two models of different sizes during the training process. The number in the figure represents Kurtosis, indicating the distribution’s tail weight and peak sharpness. High kurtosis implies heavier tails and a sharper peak, whereas low kurtosis suggests lighter tails and a flatter peak.
@@ 4.3 Degree of Emergence
While the Lipschitz–Hölder exponent $\alpha$ provides information about the nature of regularity / order and singularity, the multifractal spectrum $f(\alpha)$ offers insight into the frequency (commonly or rarely occurring pattern) and thus can help us quantify the complex and heterogeneous nature of a weighted network topology. Here we extract measures from the multifractal spectrum to quantify the regularity and heterogeneity of network structures, and use these to construct our measure for the degree of emergence.
Definition 4.5 (Irregularity Metric $\alpha_{0}$).
The exponent $\alpha_{0}$ is defined as the Lipschitz–Hölder exponent value $\alpha$ at which the multifractal spectrum $f(\alpha)$ attains its maximum, i.e.,
@@
Definition 4.6 (Heterogeneity Metric $w$).
The width $w$ of the multifractal spectrum is defined as the difference between the maximum and minimum values of the Lipschitz–Hölder exponent $\alpha$, represented by
@@
To quantify the emergence, which are critical for understanding the learning process in LLMs, we estimate the order / regularity $\alpha_{0}(t)$ and heterogeneity $w(t)$ at epoch $t$ from their corresponding network representation.
Figure 6: Multifractal spectra of Neuronal Interaction Networks (NINs) for models of varying sizes throughout the training process.
Figure 7: The comparison between our proposed metric with two metrics.
@@
This formula integrates the changes in heterogeneity and regularity over time, where $\frac{w(t)}{w(0)}$ signifies the relative change in heterogeneity, i.e., diversity in network structures, and $\log\left(\frac{\alpha_{0}(0)}{\alpha_{0}(t)}\right)$ represents the change in regularity, providing a normalized view of structural shifts. The product of these terms yields a comprehensive measure of network’s evolving complexity and organization, offering insights into the nature and degree of emergence within the system. See Section A.2.3 for a more complete explanation.
5 Experiment
@@
In this section, we describe the experiments conducted for analyzing the NINs to observe the self-organization process during training and evaluating our new metrics. Our experiments are conducted on Pythia (Biderman et al., 2023) models (except for specific requirements like different architectures). Pythia models provide checkpoints during training phases and different model scales, which is helpful for us to reveal the rules behind these scales of training step and model size.
5.1 Average weighted degree distribution
Figure 5 shows a clear distinction between the two models. During the training process, the average degree distribution for pythia 14M exhibits a dispersed trend, with lower kurtosis values (where kurtosis is commonly used to measure distribution characteristics, with higher kurtosis indicating a heavier-tailed distribution, see Section A.4), the distribution for pythia 1.4B with left shifting indicates the trend of appearing a heavy-tail distribution over multiple epochs of pertaining, which is also indicated by the increasing Kurtosis value. Please refer to Section C.4 for more information and results of Kurtosis value on pythis 14M, 160M, and 1.4B. This indicates that as the scale of the model increases, the average distance from most neurons to the neighbors in the next layer decreases, signifying the establishment of stronger interactions. It might imply that an organizational structure with a heavy-tailed connectivity pattern systematically self-organizes. Therefore, further in-depth research into the network’s connectivity structure is required to observe the self-organization process, leading us to the multifractal analysis of the network.
5.2 The Self-organization Process of the LLM Networks
@@
We analyze the dynamic LLM networks during their training process. The self-organization process, characterized by the emergence of new patterns and increasingly organized structures, can be quantitatively assessed using the $w$ and $\alpha_{0}$ metrics from Section 4.3. For each model with different sizes, NeuroMFA is applied to networks at various epochs. The curves showing the changes in the values of $\alpha_{0}$ and $w$ with epochs are provided in Fig. 14. The spectra, which vary with the epochs, exhibit distinct behaviors that reveal certain trends, as discussed below.
@@
Increasing Heterogeneity. The spectrum width $w$ measures the degree of heterogeneity of the network. Figure 6 shows that $w$ tends to broaden with more training epochs. This widening, evident from the initial spectrum to later stages, signals an increasing heterogeneity in the network’s structure. This phenomenon is largely due to the emergence of varied fractal patterns, which are intricate structures formed through neuronal interactions during training. These patterns contribute to the spectrum width’s expansion.
@@
In addition, as shown in Fig. 6, the spectra become wider with increasing epochs, suggesting an increase in the heterogeneity of the network structure. This widening is attributed to the formation of more interactions exhibiting different fractal patterns during the training process, thus broadening the spectrum.
Decreasing Irregularity. In network analysis, a leftward shift of the multifractal spectrum indicates self-organization. This shift suggests an increase in the network’s regularity, as marked by a decrease in the $\alpha$ parameter. However, Fig. 6 (a) shows for smaller model (e.g., 14M) an increasing trend of $\alpha_{0}$. As training progresses through epochs, the spectrum tends to shift rightward slightly instead of leftward. This indicates that, in the 14M model, the network’s regularity does not increase. Similarly, the 31M model displays a subtle leftward shift at the 5,000th epoch, but this trend does not continue significantly thereafter.
In contrast, for larger models (exceeding 100M parameters), a consistent leftward shift of the network spectra is observed from the 0th to the 35,000th epoch. This shift suggests an increase in regularity and a corresponding enhancement in self-organization as training progresses. After reaching the 3,5000th epoch, this leftward shift in the spectrum stabilizes across these larger models. This stabilization implies that further increases in training epochs do not significantly affect the network’s degree of regularity. This could indicate that the model training has reached a certain threshold, beyond which no additional self-organization is observed.
@@ 5.3 Quantitative Analysis of Emergence
In this section, we assess the degree of emergence $E$ for LLMs during the training process. Figure 7 shows the emergence analysis across models of varying scales (160M, 1.4B, 2.8B) and their performance for different benchmarks. We observe a positive correlation, i.e., higher levels of emergence are associated with enhanced expected model performance, as evidenced by increases in testing accuracy; of note, we acknowledge that this is restricted to these datasets and a more comprehensive evaluation should be done across more datasets (see more results across different metrics in Section C.3).
@@
Data from Lambada and PIQA tests reveal a significant increase in model accuracy during the initial training phase, up to approximately the 15,000th epoch. This trend was paralleled by a swift increase in $E$. From around the 15,000th to the 40,000th epochs, we noted a gradual enhancement in model performance, aligning with the trajectory of the degree of emergence $E$. Post this phase, models tend to stabilize, exhibiting fluctuating accuracies. These fluctuations mirrored the emergence levels, further validating our metric.
@@
Moreover, we observe significant performance oscillations in two other datasets (WSC and LogiQA, see Section C.3), suggesting their limited utility in assessing model capabilities. This underscores the limitations of relying solely on performance metrics for evaluating model’s emergence. Therefore, our approach to studying emergence focuses on analyzing the model itself, rather than merely its output or performance. This methodology provides a more comprehensive view of the emergence phenomenon in LLMs, encapsulating both the intricacies of the training process and the nuanced evolution of the model. Information about the benchmarks can be found in Appendix  B.3.
@@ 5.4 Impact Evaluations
Next, we will introduce ablation experiments to investigate the effect of various designs on the emergence metric.
Varing weight transfer functions. In Section 4.1, we choose a specific weight transfer function $f$ which satisfies the property requirements. Here, we discuss if different choices of $f$ influence the analysis results. As observed, we found $\forall w\in W,0<w\leq 1$. Based on this, we choose the following alternative functions: 1).$f(w)=\coth(|w|)$ 2). $f(w)=|\log(|w|)|$ (the $\epsilon$ part remains the same).
Full analysis results can be found in Appendix C.1.1. From these results, we can see that if $f$ satisfies the property requirement we mentioned in Section 4.1, it will not influence the analysis conclusions.
Varying tasks. Although in the area of computer vision (CV) several generative models (Rombach et al., 2022) providing personalized image generation exhibit a potential degree of intelligence, no analysis of their degree of emergence exists. Thus, we conduct an emergence analysis on these models. The experimental results can be found in Appendix C.1.2. The degree of emergence for CV tasks is lower than the NLP tasks at the same parameter size, which indicates a lower level of intelligence. This may explain why CV models have not found emergence yet.
Varying layer interactions. Our neural-based MFA approach differs significantly from traditional network analysis methods in that it accounts for the directed hierarchical structure of NIN. We initially assess the local fractal properties revealed by neurons’ neighboring mass distribution.We also explored connections across multiple layers (2 and 3 layers), but the computational load was substantial, the calculation and estimation algorithms can be found at Appendix B.1. The corresponding results are presented in Appendix C.1.3. Fractal analysis at a single layer shows more pronounced changes across different epochs, demonstrating heightened sensitivity to fractal characteristic variations. Including two or three layers complicates the analysis by introducing more nodes, which might disturb the neighboring fractal rules and obscure structural changes. This complexity underscores the advantage of our method, emphasizing its ability to discern subtle shifts in structure that multi-layer analyses may miss, thereby confirming our approach’s effectiveness in capturing fractal dynamics.

### v5

@@ 3.3 Neuron-based multifractal analysis
where $D\in\mathbb{R}$ denotes the fractal dimension characterizing the fractal properties of the interactions among a neuron and its neighbors.
To capture and distinguish the neuron-based multifractality of the NIN, we develop the Neuron-based Multifractal Analysis (NeuroMFA) framework. This method integrates the fractal features of different neurons in the NIN. The distortion factor $q\in\mathbb{R}$ is introduced to differentiate the nuances of fractal structural features. For a given neuron $v_{i}$ at layer $l$ in the neural network $G=(V,E)$, the mass probability measure $p_{l,i}(r)$ at a radius $r$ is defined as $p_{l,i}(r)=\frac{N_{l,i}(r)}{T_{l,i}},$ where $N_{l,i}(r)$ is the number of neurons within the radius $r$ from neuron $v_{i}$, and $T_{l,i}$ is the total number of neighbors of neuron $v_{i}$. To characterize the fractal properties at different scales, the partition function is defined as follows:
Partition function. The partition function $Z_{q}(r)$ for the neuron-based multifractal analysis is defined as the sum of the $q$-th power of the probability measures across all neurons in the network, given by:
$Z_{q}(r)=\sum_{l=1}^{L}\sum_{v_{i}\in L_{l}}p_{l,i}(r)^{q},$ (4)
where $q$ is the distortion factor designed to distinguish different fractal features. We provide the algorithm to obtain the partition function in Appendix Algorithm 2. By capturing the power-law relationship between the partition function $Z_{q}(r)$ and the scale $\frac{r}{d}$, we have the following observation:
@@
then the network is said to possess multifractal scale invariance (Falconer, 2014). Here, $B_{i}(\epsilon)$ represents the box of scale $\epsilon$, and $\tau(q)$ is the mass exponent.
Observation 2 (Multifractal Properties).
A log-log relationship exists between $Z_{q}(r)$ and the observational scale $r/d$, validated in Appendix Table 9:
$Z_{q}(r)\sim\left(\frac{r}{d}\right)^{\tau(q)},$ (6)
where $d$ is the maximum box radius and $\tau(q)$ is the mass exponent.
$\tau(q)$ is estimated as the slope obtained from the linear regression of $\log Z_{q}(r_{k})$ against $\log\left(\frac{r_{k}}{d}\right)$:
$\tau(q)=\frac{\sum\limits_{k=1}^{m}\left(\log Z_{q}(r_{k})-\overline{\log Z_{q}(r)}\right)\left(\log\left(\frac{r_{k}}{d}\right)-\overline{\log\left(\frac{r_{k}}{d}\right)}\right)}{\sum\limits_{k=1}^{m}\left(\log\left(\frac{r_{k}}{d}\right)-\overline{\log\left(\frac{r_{k}}{d}\right)}\right)^{2}}.$ (7)
Here, $r_{k}$ represents the distance from a neuron to its $k$-th nearest neighbor within the threshold distance $d$, with $k$ ranging from 1 to $m$, where $m$ is the total number of neighbors within the box, and $r_{m}\leq d$ is the maximum box radius. The terms $\overline{\log Z_{q}(r)}$ and $\overline{\log\left(\frac{r_{k}}{d}\right)}$ denote the mean values of $\log Z_{q}(r_{k})$ and $\log\left(\frac{r_{k}}{d}\right)$, respectively. The mass exponent $\tau(q)$ quantifies the scaling behavior of the partition function $Z_{q}(r)$ across different observational scales $r$, thereby characterizing the multifractal nature of the neural interactions within the network. The algorithm for computing $\tau$ is provided in Appendix Algorithm 3.
Multifractal spectrum. The Legendre transform enables the computation of the the Lipschitz-Hölder exponent $\alpha(q)$ and multifractal spectrum $f(\alpha)$ characterizing the multifractal structural features of the network:
@@
$f(\alpha)=q\alpha(q)-\tau(q).$ (9)
The detailed algorithm to calculate the multifractal spectrum is provided in Appendix Algorithm 4.
@@ 3.4 Degree of emergence
While the Lipschitz–Hölder exponent $\alpha$ provides information about the nature of regularity/order and singularity, the multifractal spectrum $f(\alpha)$ offers insight into the frequency (commonly or rarely occurring pattern) and thus can help us quantify the complex and heterogeneous nature of weighted network topology. Here we extract measures from the multifractal spectrum to quantify the irregularity and heterogeneity of network structures, and use these to construct our measure for the degree of emergence.
Irregularity metric $\alpha_{0}$: The exponent $\alpha_{0}$ is defined as the Lipschitz–Hölder exponent value $\alpha$ at which the multifractal spectrum $f(\alpha)$ attains its maximum, i.e.,
@@
Heterogeneity metric $w$: The width $w$ of the multifractal spectrum is defined as the difference between the maximum and minimum values of the Lipschitz–Hölder exponent $\alpha$, represented by
@@
Proposed metric. To evaluate the degree of emergence of large models based on structural dynamics of NIN, we estimate two key metrics at a given epoch $t$: the order/regularity $\alpha_{0}(t)$ and heterogeneity $w(t)$. These metrics are derived from the NIN’s evolving network representation, which provides insights into the model’s internal structure as it progresses through training.
@@
$E=\frac{w(t)}{w(0)}\log\left(\frac{\alpha_{0}(0)}{\alpha_{0}(t)}\right).$ (12)
@@
This formula integrates the changes in heterogeneity and regularity over time, providing a comprehensive measure of the system’s structural evolution. Specifically, $\frac{w(t)}{w(0)}$ signifies the relative change in heterogeneity, where values greater than $1$ indicate an increase in the diversity of interaction patterns within the network. Similarly, $\log\left(\frac{\alpha_{0}(0)}{\alpha_{0}(t)}\right)$ quantifies the shift in regularity, where values greater than 00 represent the process of the network structure becoming more organized. Together, the product of these terms offers a normalized view of the structural shifts in the network, illustrating the degree of self-organization. In this context, self-organization refers to the spontaneous development of new interaction patterns and the increasing regularity within the system, which ultimately leads to the development of skills in large models. A complete explanation is provided in Section A.2.
4 Experiments
@@
In this section, we describe the experiments conducted to analyze the NINs to observe the structural dynamics process during training and evaluate our proposed metrics. Our experiments are conducted on Pythia from 14M to 2.8B (Biderman et al., 2023) models (except for specific requirements like different architectures). Pythia models provide checkpoints during training phases and different model scales, which is helpful for us to reveal the rules behind these scales of training steps and model size. A detailed description of all the benchmarks used for comparison is provided in Appendix B.6. In the following experiments, we sample 10 networks with 64 nodes per layer and average the results to obtain precise, model-independent calculations. We set parameter $\lambda$ to 1 and $\gamma$ to 5 in Eq. 1 to achieve an appropriate distance between neurons. The variance of the calculations shown in  Section D.1 demonstrates the reliability of our sampling method.
4.1 Self-organization process of the NIN in LLMs
@@
We analyze the dynamic LLM networks during their training process. The self-organization process, characterized by the emergence of new patterns and increasingly organized structures, can be quantitatively assessed using the $w$ and $\alpha_{0}$ metrics from Section 3.4. For each model with different sizes, NeuroMFA is applied to networks at various epochs. The curves showing the changes in the values of $\alpha_{0}$ and $w$ with epochs are provided in Appendix Fig. 19. The spectra, which vary with the epochs, exhibit distinct behaviors that reveal certain trends, as discussed below.
Figure 5: Multifractal spectra of Neuronal Interaction Networks (NINs) for models with different sizes (14M-2.8B) throughout the training process. The gradient of spectral lines, shifting from blue to red, represents the multifractal spectra of the same model at different epochs, with blue indicating early-stage training and red signifying later stages.
@@
Increasing heterogeneity. The spectrum width $w$ measures the degree of heterogeneity of the fractal interaction patterns within LLMs. Figure 5 shows that $w$ tends to broaden with more training epochs. This widening, evident from the initial spectrum to later stages, signals an increasing heterogeneity in the network’s structure. This phenomenon is largely due to the emergence of varied fractal patterns, which are intricate structures formed through neuronal interactions during training. These patterns contribute to the spectrum width’s expansion.
@@
Decreasing irregularity. The leftward shift of the multifractal spectrum indicates self-organization. This shift suggests an increase in the network’s regularity, as marked by a decrease in the $\alpha$ parameter (metric of irregularity). However, Fig. 5 (a) shows for the smaller model (e.g., 14M) an increasing trend of $\alpha_{0}$. As training progresses through epochs, the spectrum tends to shift rightward slightly instead of leftward. This indicates that, in the 14M model, the network’s regularity does not increase. Similarly, the 31M model displays a subtle leftward shift at the 5,000th epoch, but this trend does not continue significantly thereafter.
In contrast, for larger models (exceeding 100M parameters), a consistent leftward shift of the network spectra is observed from the 0th to the 35,000th epoch. This shift suggests an increase in regularity and a corresponding enhancement in self-organization as training progresses. After reaching the 3,5000th epoch, this leftward shift in the spectrum stabilizes across these larger models. This stabilization implies that further increases in training epochs do not significantly affect the network’s degree of regularity. This could indicate that the model training has reached a certain threshold, beyond which no additional self-organization is observed. Notably, since the degree of emergence is fundamentally tied to self-organization phenomena, we did not compute this metric for the 14M and 31M models in the following experiments.
@@ 4.2 Quantitative analysis of emergence
Figure 6: Comparison of the proposed emergence metric $E$ (blue line) with two established metrics (red line). The strong correlation between our proposed metric and existing metrics demonstrates the validity of our approach in quantifying emergent properties.
@@
In this section, we assess the degree of emergence $E$ for LLMs during the training process. Figure 6 shows the emergence analysis across models of varying scales (160M, 1.4B, 2.8B) and their performance for different benchmarks. We observe a positive correlation, i.e., higher levels of emergence are associated with enhanced expected model performance, as evidenced by increases in testing accuracy; of note, we acknowledge that this is restricted to these datasets and a more comprehensive evaluation should be done across more datasets (see more results across different metrics in Section D.2).
Figure 7: Radar charts showing the performance of (a) 1B and (b) 1.4B models across cognitive and reasoning benchmarks over different training epochs. A: LAMBADA dataset challenges, B: science exam questions, C: Physical Interaction QA, D: AI2 Reasoning Challenge - Easy.
@@
Data from Lambada and PIQA tests reveal a significant increase in model accuracy during the initial training phase, up to approximately the 15,000th epoch. This trend was paralleled by a swift increase in $E$. From around the 15,000th to the 40,000th epochs, we note a gradual enhancement in model performance, aligning with the trajectory of the degree of emergence $E$. Post this phase, models tend to stabilize, exhibiting fluctuating accuracies. These fluctuations mirror the emergence levels, further validating our metric.
@@
Moreover, we observe significant performance oscillations in two other datasets (WSC and LogiQA, see Section D.2), suggesting their limited utility in assessing model capabilities. This underscores the limitations of relying solely on performance metrics for evaluating model’s emergent abilities. Therefore, our approach to studying emergence focuses on analyzing the model itself, rather than merely its output or performance. This methodology provides a more comprehensive view of the emergence phenomenon in LLMs, encapsulating both the intricacies of the training process and the nuanced evolution of the model. Information about the benchmarks can be found in Appendix  B.6.
Degree of emergence and model capabilities. Fig. 7 compares our degree of emergence metric with established LLM benchmarks: Lambada, SciQ, PIQA, and ARC-easy. These benchmarks assess different aspects of LLM emergent abilities: Lambada evaluates contextual comprehension, testing the model’s ability to understand long-range dependencies in text. SciQ measures scientific reasoning, challenging models to apply scientific concepts beyond mere fact recall. PIQA assesses physical commonsense, requiring models to reason about everyday physical interactions. ARC-easy tests basic science knowledge and fundamental reasoning across various disciplines. emeTogether, these metrics provide a comprehensive assessment of LLM capabilities, covering language understanding, scientific reasoning, commonsense knowledge, and basic problem-solving skills. We analyze Pythia 1B and Pythia 1.4B models across different training epochs. The radar charts demonstrate that our proposed degree of emergence aligns closely with these established metrics as training progresses.
Figure 8: Degree of Emergence vs. Model Size. Avg. Performance means the average score of all the metrics. All metric scores have been normalized between 0 and 1 for comparison.
Our metric reflects the gradual enhancement of LLM’s overall capabilities as the training epoch increases, mirroring the trends observed in other LLM benchmarks. This correlation suggests that our degree of emergence, derived from the NeuroMFA method, effectively captures the comprehensive development of LLMs over time. The consistent behavior between our structure-based metric and the diverse set of established performance benchmarks throughout the training process validates the NeuroMFA approach. It also demonstrates the feasibility of assessing LLM capabilities from a structural perspective, offering a new pathway to evaluate the comprehensive abilities of language models based on their internal organization rather than just output performance.
Degree of emergence vs. model size. We evaluate the degree of emergence across models of varying sizes by comparing their performance on different benchmarks, specifically PIQA, LAMBADA, and ARC Challenge, along with the average score across all evaluation metrics. As illustrated in Fig. 8, at an early stage of training (512 epochs), all models, regardless of size, show a low degree of emergence, exhibit a low degree of emergence, reflecting limited emergent abilities across the benchmarks. However, at a later stage of training (53,000 epochs), the degree of emergence increases logarithmically with model size, suggesting that larger models demonstrate more pronounced emergent abilities across a range of tasks.
@@ 4.3 Impact evaluations
We applied NeuroMFA to analyze varying architectures (e.g., diffusion models, CNNs) and training data (e.g., contaminated vs. clean datasets, Pythia-standard vs. Pythia-deduped). The results, detailed in Appendix C, highlight differences in emergent abilities across architectures and the impact of training data on self-organization and emergence.